# Quantum walks and entanglement in cavity networks

### v1

@@ Abstract
For harnessing the full potential of quantum phenomena, light-matter interfaces and complexly connected quantum networks are required, relying on the joint quantum operation of different physical platforms. In this work, we analyze the quantum properties of multipartite quantum systems, consisting of an arbitrarily large collection of optical cavities with two-level atoms. In particular, we explore quantum walks in such systems and determine the resulting entanglement. Realistic imperfections are included in the model as optical losses and spontaneous decays of atoms. The topology of torus and the non-orientable Möbius strip serve as examples of complex networks we consider, demonstrating the versatility of our approach and resulting in interesting quantum dynamics and interference effects for quantum simulation applications.
@@ I Introduction
Because of the aforementioned relevance, quantum walks have been implemented in a large variety of physical systems. In quantum optics, for example, this is usually achieved via various linear optical networks [24, 25, 26, 27, 28, 29, 30]. For instance, input nonclassicality is converted to output entanglement in such scenarios via simple beam splitters [31, 32, 33]. Other optical schemes are based, for example, on photonic lattices [34, 35, 36, 37, 38]. Moreover, quantum simulators with a nonlinear quantum walk evolution have been studied [39, 40, 41, 42, 43, 44, 45], and imperfect quantum walks on graphs with decoherence have been investigated [46, 47]. Furthermore, quantum walks have been realized with single optically trapped atoms [48] and in ion traps [49, 50].
@@
More generally, trapped atoms and ions interacting with quantized light fields are promising candidates for the realization of hybrid quantum information processing [51, 52, 53, 54]. The combination of long-lived atomic states and light fields can be highly beneficial in quantum networking and for distributed quantum computation [55, 56, 57]. Moreover, various achievements in cavity QED and in tapped-ion techniques have rendered it possible to experimentally generate pairs of entangled atoms [58], to create entangled states of several atoms [59], and to maintain robust entanglement of macroscopic ensembles of atoms [60]. Light-matter interfaces also naturally provide a link between the continuous-variable entanglement of photons [61] and cold atoms [62].
@@
In the present paper, we combine entanglement theory, quantum simulation, and cavity QED to study the generation and distribution of entanglement in a network. In view of the widespread applications of cavity-assisted single-photon sources, we analyze composite quantum systems consisting of atom-cavity nodes that are connected in a cascaded configuration [63, 64, 65]. As found for smaller cascades [66, 67, 68], one of the advantages of using such configurations is that the open-system evolution itself creates the entanglement [69, 70, 71, 72], even extending to other scenarios, such as optomechanical systems [73]. Here, we consider an arbitrary network of cavities that are optically coupled, including realistic imperfections. Initializing the system with a single excitation, we explore how this excitation propagates between quantized radiation fields and the matter and is simultaneously spread across the network. This implements a quantum walk that can be compared with the analogous classical random walk. Moreover, we analyze the multipartite entanglement dynamics that is established in the networks between the different nodes, as well as the light-matter entanglement, by applying tailored multipartite entanglement witnesses. Elaborate network configurations as non-trivial graphs are characterized.
@@
The paper is organized as follows. In Sec. II, the master equation describing the dynamics of a single atom-cavity system is reviewed. In Sec. III, the master equation describing the dynamics of arbitrary multi-cavity system is formulated, and the evolution is solved analytically by means of the quantum trajectory method. In Sec. IV, we study different network geometries, specifically a network with a toroidal structure and one with a Möbious-strip configuration. We further analyze how an initial excitation propagates, providing a light-matter-interfacing quantum walk. In Sec. V, the entanglement between the different constituents of the full quantum system is analyzed, including the impact of attenuations. Finally, concluding remarks are offered in Sec. VI.
@@ II Damped Light-Matter Dynamics
Before considering an entire network, we analyze a single building block that is used in the next section as a node to construct the cascaded network. Specifically, we consider a high-$Q$ optical cavity, supporting one monochromatic mode, $\omega^{(C)}$, that is coupled to a two-level atomic transition, $\omega^{(A)}$; see Fig. 1. The detuning is given by the difference of both frequencies. The included imperfections are: the cavity mode is damped by losses through partially transmitting cavity mirrors; in addition to wanted in- and our-coupling, the atom can spontaneously emit out of the side of the cavity; and photons may be absorbed and scattered by cavity mirrors [74].
@@
with $g$ denoting the atom-cavity coupling constant. Note that an inter-cavity interaction, coupling the incoming and outgoing fields, is not present here but is introduced in the next section. It is also worth recalling that solving the master equation, as commonly written in the Lindblad form [78], can be achieved through the quantum trajectory approach [76, 77, 75].
@@ III.1 Full multipartite system
where $\mathcal{J}$ denotes the set of all jump operators (aka Lindblad operators). The full Hamiltonian is given by
@@
with the coupling $\kappa_{l,m}$ between cavities $l$ and $m$, obeying the conditions $\kappa_{l,m}=\kappa_{m,l}^{\ast}$ and $\kappa_{l,l}=0$ (no self-coupling). For coupling multiple cavities, other than in Fig. 1, one can, for example, utilize a ring-cavity configuration with multiple mirrors, each of which providing additional input and output ports for coupling [79, 80].
@@
For the proof of concept, we restrict ourselves to a single excitation (likewise, one quantum walker) throughout this work. For the purpose of quantum walks, the $k$th mode acts as the position-like degree of freedom and the state of the two-level atom functions as the quantum coin. For convenience, we further define the following state vectors:
@@
Here, $|g\rangle$ is the state of the network in which all the atoms are found in their ground state and all the cavity modes are void of photons, vacuum state. In addition, $|1_{k}^{(C)}\rangle$ says that the $k$th cavity is occupied with one photon, and $|1_{k}^{(A)}\rangle$ indicates that the $k$th two-level atom is excited.
@@ III.2 Analytic solution
Figure 2: Random walks on graphs. We consider different, interesting geometries, a torus (a1) as a nontrivial topology and a Möbius strip (b1) as a non-orientable manifold, where vertices are depicted as black circles and connecting edges as black lines. Vertices are located as pairs of positions across and around the torus and Möbius strip. For later comparisons with the quantum walks, classical random walks are implemented in (a2) and (b2) for both graphs [81]. The walker is initialized at vertex $(3,3)$ and randomly proceeds to neighboring positions. With propagation in time, we see a diffusive behavior, eventually leading to a uniform distribution. Here we made the particular choice of 25 vertices, i.e., $N_{1}=N_{2}=5$. In panels (a3) and (b3), the probability distribution of an intermediate time is shown for the two geometries. For the quantum walk, cavities resemble the vertices of the graphs and the coupling between cavities is indicated by edges.
@@ IV Continuous-in-time quantum walks on different geometries
In this section, we study different network geometries by analyzing how the initially localized excitation propagates through the network. By exploring specific examples of a non-trivial quantum walk dynamics, we demonstrate the flexibility of the model. The two cases under study pertain to a network on a closed surface with a toroidal structure and a network on a two-dimensional non-orientable surface in a Möbious-strip configuration; see also Fig. 2, including the classical random-walk evolution.
@@ IV.1 Considered scenarios and random walk
The characterization of the chosen geometry is realized by specifying the particular interaction Hamiltonian, Eq. (6), that describes the interaction between the different cavities. This requires to specify the matrix $\bm{K}=[\kappa_{i,j}]_{i,j}$ in Eq. (14). For investigating different geometries, we consider a network as a graph wherein the vertices, i.e., nodes, are the cavities together with the two-level atom, and edges of the graph define the network connectivity. By applying Eq. (15), the dynamics of the system in the chosen network geometry is then fully determined.
In the scenarios described above, the adjacency matrix $\bm{A}$, well-known from graph theory [82], is proportional to the coupling matrix. Further, we assume a uniform $\kappa$ coupling, where $\kappa>0$, between connected vertices and undirected channels, $\bm{A}=\bm{A}^{\mathrm{T}}$. Therefore, we can write $\bm{K}$ as
@@
The first geometry we consider is a closed surface with a toroidal structure, Fig. 2 (a1). Topologically speaking, the torus has a non-trivial genus one. A graph on a torus is characterized by nodes $i=(i_{1},i_{2})$, a two-dimensional grid with $1\leq i_{1},j_{1}\leq N_{1}$ and $1\leq i_{2},j_{2}\leq N_{2}$. In addition, a periodic boundary condition is imposed on the nodes in both components separately. Therefore, the entries of the adjacency matrix $\bm{A}$ can be expressed as
@@
The second example pertains to a two-dimensional graph configuration on a Möbious strip, cf. Fig. 2 (b1). In contrast to the previous graph, this strip corresponds to a non-orientable surface. For this graph, the adjacency matrix $\bm{A}$ is determined by the entries
@@
Figure 3: Dynamics of the quantum walk for a network with toroidal structure (top row) and a Möbious strip (bottom row). The probability of finding the walker at a certain position as a cavity excitation and atom excitation is shown in the first and second columns, respectively. Accordingly, the evolution of the cavity probability $|\alpha_{k=(k_{1},k_{2})}(t)|^{2}$ on each configuration is shown in (a1) and (b1), and the evolution of the atomic probability $|\beta_{k=(k_{1},k_{2})}(t)|^{2}$, in (a2) and (b2) for discrete time steps $t\to t+dt$. In the last two columns, the probability distribution of an intermediate time is depicted. In particular, for a chosen time step $t=40\,dt$, the cavity probability $|\alpha_{k=(k_{1},k_{2})}|^{2}$ is shown in (a3) and (b3), and the atomic probability $|\beta_{k=(k_{1},k_{2})}|^{2}$ in (a4) and (b4). For the strong-coupling regime illustrated here, the considered parameters in units of the coupling $\kappa$ are $\Gamma^{(C)}/\kappa=\Gamma^{(A)}/\kappa=0$, $\omega^{(A)}/\kappa=1$, $\omega^{(C)}/\kappa=0.9$ and $g/\kappa=10^{5}$, while defining $dt=1/\kappa$ in the numerical analysis.
@@ IV.2 Quantum evolution
After having introduced the geometries of interest and the dynamics of classical random walks, we now analyze the quantum evolution described by our model. For this purpose, we compute the dynamics of a quantum walk according to Eq. (15) for both geometries. For the sake of comparison, we analyze the same scenario presented in Fig. 2; that is, a network of $N_{1}=N_{2}=5$ cavities and, as initial condition, a single excitation of the cavity field located at vertex $(3,3)$ and no excitation of atoms, $|\bar{\psi}(t_{0}=0)\rangle=|1^{(C)}_{k}\rangle$ with $k=(3,3)$.
The results of our computation are presented in Fig. 3, where the top row corresponds to the torus and the bottom one to the Möbius strip. The evolution of the probability is given as a timeline in the first two columns. The probability of finding the excitation as a photon across the nodes, $|\alpha_{k}|^{2}$, appears in panels (a1) and (b1), whereas the probability of finding an excited atom, $|\beta_{k}|^{2}$, is presented in panels (a2) and (b2). Note that occupation probability significantly alternates between light and matter degrees of freedom with a frequency modulated by $\bm{\Omega}$, cf. Eqs. (16) and (17). The most striking signature of the quantum behavior is the interference between the probability amplitudes, which gives rise to rich interference patterns in the probability distribution.
This interference, which can be already seen in the four timelines, is even more clear if we inspect one specific time step. This is shown in Fig. 3 in panels (a3) and (b3) for photon probabilities and in panels (a4) and (b4) for excited-atom probabilities. By comparing the same time step in the torus and Möbius strip, one can see the difference between the structures of the interference, which is not observable in the random walk analysis in Fig. 2. In particular, there is a clear asymmetry between vertices’ position component $1$ and $2$ for the probability distribution in the Möbius strip. By contrast, position components $1$ and $2$ are interchangeable for the torus in this particular scenario. Importantly, we emphasize that all interference patterns noticeably contrast with the interference-free distribution in the random-walk scenario; compare panels (a3) and (b3) in Fig. 2 with panels (a3), (a4), (b3), and (b4) in Fig. 3.
For comparison with the random walk, we assumed lossless cases here; the impact of quantum jumps is discussed in the next section together with characterizing the systems’ entanglement. Furthermore, one way to extend our model is to introduce an optical inter-cavity field that couples to the internal field of connected cavities studied here. Losses of inter-cavity fields act as an additional decoherence mechanism, leading to a transition from strong interference to classical diffusion as observed in a random walk [47, 48, 69].
@@ V Multipartite entanglement evolution
In the previous section, we focused on the classical and quantum probabilities to find the walker in a certain configuration. However, this not necessarily proves that quantum correlations are generated through the evolution. Therefore, we now provide a detailed entanglement analysis of the systems under consideration.

### v2

@@ Abstract
For harnessing the full potential of quantum phenomena, light-matter interfaces and complexly connected quantum networks are required, relying on the joint quantum operation of different physical platforms. In this work, we analyze the quantum properties of multipartite quantum systems, consisting of an arbitrarily large collection of optical cavities with two-level atoms. In particular, we explore quantum walks in such systems and determine the resulting entanglement. Realistic imperfections are included in the model as optical losses and spontaneous decays of atoms. The topology of torus and the nonorientable Möbius strip serve as examples of complex networks that we consider, demonstrating the versatility of our approach and resulting in interesting quantum dynamics and interference effects for quantum simulation applications.
@@ I Introduction
Because of the aforementioned relevance, quantum walks have been implemented in a large variety of physical systems. In quantum optics, for example, this is usually achieved via various linear optical networks [24, 25, 26, 27, 28, 29, 30]. For instance, input nonclassicality is converted to output entanglement in such scenarios via simple beam splitters [31, 32, 33]. Other optical schemes are based, for example, on photonic lattices [34, 35, 36, 37, 38]. Moreover, quantum simulators with a nonlinear quantum-walk evolution have been studied [39, 40, 41, 42, 43, 44, 45], and imperfect quantum walks on graphs with decoherence have been investigated [46, 47]. Furthermore, quantum walks have been realized with single optically trapped atoms [48] and in ion traps [49, 50]. Another promising quantum-walk platform is optomechanical and electromechanical systems [51, 52, 53].
@@
In general, trapped atoms and ions interacting with quantized light fields are well-suited candidates for the realization of hybrid quantum information processing [54, 55, 56, 57]. The combination of long-lived atomic states and light fields can be highly beneficial in quantum networking and for distributed quantum computation [58, 59, 60]. Moreover, various achievements in cavity QED and in tapped-ion techniques have rendered it possible to experimentally generate pairs of entangled atoms [61], to create entangled states of several atoms [62], and to maintain robust entanglement of macroscopic ensembles of atoms [63]. Light-matter interfaces also naturally provide a link between the continuous-variable entanglement of photons [64] and cold atoms [65].
@@
In the present contribution, we combine entanglement theory, quantum simulation, and cavity QED to study the generation and distribution of entanglement in networks. In view of the widespread applications of cavity-assisted single-photon sources, we analyze composite quantum systems consisting of atom-cavity nodes that are connected in a cascaded configuration [66, 67, 68]. As found for smaller cascades [69, 70, 71], one of the advantages of using such configurations is that the open-system evolution itself creates the entanglement [72, 73, 74, 75], even extending to other scenarios, such as optomechanical systems [76]. Here, we consider an arbitrary network of cavities that are optically coupled, including realistic imperfections. Initializing the system with a single excitation, we explore how this excitation propagates between quantized radiation fields and the matter and is simultaneously spread across the network. This implements a quantum walk that can be compared with the analogous classical random walk. Moreover, we analyze the multipartite entanglement dynamics that is established in the networks between the different nodes, as well as the light-matter entanglement, by applying tailored multipartite entanglement witnesses. Elaborate network configurations as nontrivial graphs are characterized.
@@
The paper is organized as follows. In Sec. II, the master equation describing the dynamics of a single atom-cavity system is reviewed. In Sec. III, the master equation describing the dynamics of an arbitrary multi-cavity system is formulated, and the evolution is solved analytically by means of the quantum trajectory method. In Sec. IV, we study different network geometries, specifically a network with a toroidal structure and one with a Möbius-strip configuration. We further analyze how an initial excitation propagates, providing a light-matter-interfacing quantum walk. In Sec. V, the entanglement between the different constituents of the full quantum system is analyzed, including the impact of attenuations. Finally, concluding remarks are offered in Sec. VI.
@@ II Damped Light-Matter Dynamics
Before considering an entire network, we analyze a single building block that is used in the next section as a node to construct the cascaded network. Specifically, we consider a high-$Q$ optical cavity, supporting one monochromatic mode, $\omega^{(C)}$, that is coupled to a two-level atomic transition $\omega^{(A)}$; see Fig. 1. The detuning is given by the difference of both frequencies. The included imperfections are as follows: the cavity mode is damped by losses through partially transmitting cavity mirrors; in addition to wanted in- and out-coupling, the atom can spontaneously emit out of the side of the cavity; and photons may be absorbed and scattered by cavity mirrors [77].
@@
with $g$ denoting the atom-cavity coupling constant. Note that an intercavity interaction, coupling the incoming and outgoing fields, is not present here but is introduced in the next section. It is also worth recalling that solving the master equation, as commonly written in the Lindblad form [81], can be achieved through the quantum trajectory approach [79, 80, 78].
@@ III.1 Full multipartite system
where $\mathcal{J}$ denotes the set of all jump operators (also known as Lindblad operators). The full Hamiltonian is given by
@@
with the coupling $\kappa_{l,m}$ between cavities $l$ and $m$, obeying the conditions $\kappa_{l,m}=\kappa_{m,l}^{\ast}$ and $\kappa_{l,l}=0$ (no self-coupling). For coupling multiple cavities, other than in Fig. 1, one can, for example, utilize a ring-cavity configuration with multiple mirrors, each of which provides additional input and output ports for coupling [82, 83].
@@
As a proof of concept, we restrict ourselves to a single excitation (likewise, one quantum walker) throughout this work. For the purpose of quantum walks, the $k$th mode acts as the position-like degree of freedom, and the state of the two-level atom functions as the quantum coin. For convenience, we further define the following state vectors:
@@
Here, $|g\rangle$ is the state of the network in which all the atoms are found in their ground state and all the cavity modes are void of photons, i.e., avacuum state. In addition, $|1_{k}^{(C)}\rangle$ says that the $k$th cavity is occupied with one photon, and $|1_{k}^{(A)}\rangle$ indicates that the $k$th two-level atom is excited.
@@ III.2 Analytic solution
Figure 2: Random walks on graphs. We consider different, interesting geometries, i.e., (a1) a torus as a nontrivial topology and (b1) a Möbius strip as a nonorientable manifold, where vertices are depicted as black circles and connecting edges as black lines. Vertices are located as pairs of positions across and around the torus and Möbius strip. For later comparisons with the quantum walks, classical random walks are implemented in (a2) and (b2) for both graphs [84]. The walker is initialized at vertex $(3,3)$ and randomly proceeds to neighboring positions. With propagation in time, we see a diffusive behavior, eventually leading to a uniform distribution. Here we made the particular choice of 25 vertices, i.e., $N_{1}=N_{2}=5$. In panels (a3) and (b3), the probability distribution of an intermediate time is shown for the two geometries. For the quantum walk, cavities resemble the vertices of the graphs and the coupling between cavities is indicated by edges.
@@ IV Continuous-in-time quantum walks on different geometries
In this section, we study different network geometries by analyzing how the initially localized excitation propagates through the network. By exploring specific examples of a nontrivial quantum-walk dynamics, we demonstrate the flexibility of the model. The two cases under study pertain to a network on a closed surface with a toroidal structure and a network on a two-dimensional nonorientable surface in a Möbius-strip configuration; see, also, Fig. 2, including the classical random-walk evolution.
@@ IV.1 Considered scenarios and random walk
The characterization of the chosen geometry is realized by specifying the particular interaction Hamiltonian, given by Eq. (6), that describes the interaction between the different cavities. This requires one to specify the matrix $\bm{K}=[\kappa_{i,j}]_{i,j}$ in Eq. (14). To investigate different geometries, we consider a network as a graph wherein the vertices, i.e., nodes, are the cavities together with the two-level atom, and edges of the graph define the network connectivity. By applying Eq. (15), the dynamics of the system in the chosen network geometry is then fully determined.
In the scenarios described above, the adjacency matrix $\bm{A}$, well known from graph theory [85], is proportional to the coupling matrix. Further, we assume a uniform $\kappa$ coupling, where $\kappa>0$, between connected vertices and undirected channels, $\bm{A}=\bm{A}^{\mathrm{T}}$. Therefore, we can write $\bm{K}$ as
@@
The first geometry we consider is a closed surface with a toroidal structure; see Fig. 2(a1). Topologically speaking, the torus has a nontrivial genus of one. A graph on a torus is characterized by nodes $i=(i_{1},i_{2})$, and a two-dimensional grid with $1\leq i_{1},j_{1}\leq N_{1}$ and $1\leq i_{2},j_{2}\leq N_{2}$. In addition, a periodic boundary condition is separately imposed on the nodes in both components. Therefore, the entries of the adjacency matrix $\bm{A}$ can be expressed as
@@
The second example pertains to a two-dimensional graph configuration on a Möbius strip; cf. Fig. 2(b1). In contrast to the previous graph, this strip corresponds to a nonorientable surface. For this graph, the adjacency matrix $\bm{A}$ is determined by the entries
@@
Figure 3: Dynamics of the quantum walk for a network with toroidal structure (top row) and a Möbius strip (bottom row). The probability of finding the walker at a certain position as a cavity excitation and atom excitation is shown in the first and second columns, respectively. Accordingly, the evolution of the cavity probability $|\alpha_{k=(k_{1},k_{2})}(t)|^{2}$ on each configuration is shown in (a1) and (b1), and the evolution of the atomic probability $|\beta_{k=(k_{1},k_{2})}(t)|^{2}$ is shown in (a2) and (b2) for discrete time steps $t\to t+dt$. In the last two columns, the probability distribution of an intermediate time is depicted. In particular, for a chosen time step $t=40\,dt$, the cavity probability $|\alpha_{k=(k_{1},k_{2})}|^{2}$ is shown in (a3) and (b3), and the atomic probability $|\beta_{k=(k_{1},k_{2})}|^{2}$ is shown in (a4) and (b4). For the strong-coupling regime illustrated here, the considered parameters in units of the coupling $\kappa$ are $\Gamma^{(C)}/\kappa=\Gamma^{(A)}/\kappa=0$, $\omega^{(A)}/\kappa=1$, $\omega^{(C)}/\kappa=0.9$, and $g/\kappa=10^{5}$, while defining $dt=1/\kappa$ in the numerical analysis.
@@ IV.2 Quantum evolution
After having introduced the geometries of interest and the dynamics of classical random walks, we now analyze the quantum evolution described by our model. For this purpose, we compute the dynamics of a quantum walk according to Eq. (15) for both geometries. For the sake of comparison, we analyze the same scenario presented in Fig. 2; that is, a network of $N_{1}=N_{2}=5$ cavities and, as the initial condition, a single excitation of the cavity field located at vertex $(3,3)$ and no excitation of atoms, $|\bar{\psi}(t_{0}=0)\rangle=|1^{(C)}_{k}\rangle$ with $k=(3,3)$.
The results of our computation are presented in Fig. 3, where the top row corresponds to the torus and the bottom one to the Möbius strip. The evolution of the probability is given as a timeline in the first two columns. The probability of finding the excitation as a photon across the nodes, $|\alpha_{k}|^{2}$, appears in Figs. 3(a1) and (b1), whereas the probability of finding an excited atom, $|\beta_{k}|^{2}$, is presented in Figs.(a2) and (b2). Note that occupation probability significantly alternates between light and matter degrees of freedom with a frequency modulated by $\bm{\Omega}$; cf. Eqs. (16) and (17). The most striking signature of the quantum behavior is the interference between the probability amplitudes, which gives rise to rich interference patterns in the probability distribution.
This interference, which can already be seen in the four timelines, is even more clear if we inspect one specific time step. This is shown in Figs. 3(a3) and 3(b3) for photon probabilities and in Figs. 3(a4) and 3(b4) for excited-atom probabilities. By comparing the same time step in the torus and Möbius strip, one can see the difference between the structures of the interference, which is not observable in the random walk analysis in Fig. 2. In particular, there is a clear asymmetry between the vertices’ position component $1$ and $2$ for the probability distribution in the Möbius strip. By contrast, position components $1$ and $2$ appear interchangeable for the torus in this particular scenario. Importantly, we emphasize that all interference patterns noticeably contrast with the interference-free distribution in the random-walk scenario; compare Figs. 2(a3) and 2(b3) with Figs. 3(a3), 3(a4), 3(b3), and 3(b4).
For the sought-after comparison with the random walk, we assumed lossless cases here; the impact of quantum jumps is discussed in the next section together with characterizing the systems’ entanglement. Furthermore, one way to extend our model is to introduce an optical intercavity field that couples to the internal field of connected cavities studied here. Losses of intercavity fields act as an additional decoherence mechanism, leading to a transition from strong interference to classical diffusion as observed in a random walk [47, 48, 72].
@@ V Multipartite entanglement evolution
In the previous section, we focused on the classical and quantum probabilities to find the walker in a certain configuration. However, this does not necessarily prove that quantum correlations are generated through the evolution. Therefore, we now provide a detailed entanglement analysis of the systems under consideration.